OpenStudy (anonymous):

What is the result of (6.2 × 10^4) x (3.3 × 10^2) expressed in scientific notation

OpenStudy (anonymous):

You have powers of 10 and multiplication so you can multiply your numbers and add up your exponents of powers of 10.

OpenStudy (anonymous):

\[(6.2 \times 3.3) \times 10 ^ {4+2}\]

OpenStudy (anonymous):

Do the multiplication and the addition as well as when you multiply out the 6.2 and 3.3 you will need to convert that number into another additional power 10.

OpenStudy (anonymous):

\[6.2 \times 10^4 = 62000\]